Pam is 3 years older than her sister, Jan. The sum of their ages is 99. Find their ages.
Answer by ikleyn(52787)

.

Jan + Pam       = 99


Jan + (Jan + 3) = 99


2*Jan + 3 = 99


2*Jan = 99 - 3 = 96.


Jan = 96/2 = 48.      ANSWER


Pam = 48+3 = 51.      ANSWER